# egoPPG: Heart Rate Estimation from Eye-Tracking Cameras in Egocentric Systems to Benefit Downstream Vision Tasks

### v1

@@ 10 Related datasets
Tab. 6 gives a comprehensive comparison of the dataset size and activities of related remote photoplethysmography (rPPG) datasets. In terms of hours of recordings and recorded frames, egoPPG-DB is the third largest dataset. Furthermore, we see that all comparable rPPG datasets only include activities with very little motion and heart rate (HR) changes such as watching videos, head rotations or talking. In contrast, egoPPG-DB features a wide variety of challenging everyday activities, such as kitchen work, dancing riding and exercise bike, which induce significant motion artifacts and HR changes.
Dataset Part. Frames Hours Tasks
PURE [78] 10 110 K 1 Resting, talking, small head movements
MAHNOB-HCI [77] 27 2.6 M 12 Watching videos
MMPD [80] 33 1.2 M 11 Resting, head rotation, selfie videos
UBFC-rPPG [7] 43 150 K 1.5 Gaming on a computer
UBFC-PHYS [70] 56 2.4 M 19 Resting, Trier Social Stress Test
Table 6: Summary of existing datasets for rPPG.
@@ 11 Excluded tasks
For all participants and activities, we checked the mean absolute error (MAE) between the predicted HR from our custom contact PPG sensor on the nose and the gold standard ECG from the chest belt. We excluded all tasks with an MAE over 3.0 beats per minute (bpm), which can happen, for example, when the PPG sensor loses alignment with the angular artery due to movement. In this way, we ensured that the photoplethysmography (PPG) signal signal from the nose, which we used as the target signal to train our model, is highly accurate. As a result, we had to exclude 20 out of the 150 tasks (13%), which we list in Tab. 7. We can see that this applied only to tasks with more motion (dancing, exercise bike, and walking). Since the participants had to walk multiple stairs throughout the data recording, this mostly happened during walking.
@@ 12 Detailed description of activities
Walking Walk back to the start Walk back to the start either up the stairs or using the elevator.
@@ 13 Data recording
14 Initial signal verification
15 Hyperparameter optimization
In Tab. 9, we show the results of EgoPulseFormer when training our model with different hyperparameters such as different image sizes, input window sizes, or target signals.
Hyperparameter Configuration MAE RMSE MAPE r
Image preprocessing Raw 16.33 20.04 20.93 0.31
Standardized 14.44 18.21 18.66 0.46
Difference 8.99 12.09 10.96 0.80
Standardized difference 8.82 12.03 10.82 0.81
Image size $24\times 64$ 9.73 12.69 12.12 0.77
$\boldsymbol{48\times 128}$ 8.82 12.03 10.82 0.81
$96\times 256$ 10.27 13.10 12.46 0.79
Window size 64 9.95 12.82 11.93 0.80
128 8.82 12.03 10.82 0.81
256 9.71 13.02 11.41 0.78
Loss function Negative Pearson [92] 10.62 13.57 13.16 0.81
MAE 9.71 13.02 11.41 0.78
MSE 8.82 12.03 10.82 0.81
Target signal Heart rate 10.99 14.03 13.49 0.77
Nose PPG 8.82 12.03 10.82 0.81
Augmentation Without 10.57 14.19 12.78 0.70
With 8.82 12.03 10.82 0.81
Table 9: Results of our model on the egoPPG-DB dataset when training with different hyperparameters.

### v2

@@ 10 Related datasets
Tab. 10 gives a comparison of the dataset size and activities of some related remote photoplethysmography (rPPG) datasets. In terms of hours of recordings and recorded frames, egoPPG-DB is among the largest dataset. Furthermore, we see that all comparable rPPG datasets only include activities with very little motion and heart rate (HR) changes such as watching videos, head rotations or talking. In contrast, egoPPG-DB features a wide variety of challenging everyday activities, such as kitchen work, dancing and riding an exercise bike, which induce significant motion artifacts and HR changes.
@@ 11 Excluded tasks
For all participants and activities, we checked the mean absolute error (MAE) between the predicted HR from our custom contact PPG sensor on the nose and the gold standard ECG from the chest belt. We excluded all tasks with an MAE over 3.0 beats per minute (bpm), which can happen, for example, when the PPG sensor loses alignment with the angular artery due to movement. In this way, we ensured that the photoplethysmography (PPG) signal from the nose, which we used as the target signal to train our model, is highly accurate. As a result, we had to exclude 20 out of the 150 tasks (13%), which we list in Tab. 6. We can see that this applied only to tasks with more motion (dancing, exercise bike, and walking). Since the participants had to walk multiple stairs throughout the data recording, this mostly happened during walking.
@@ 12 Detailed description of activities
13 Data recording
14 Initial signal verification
15 Variance of results
In Fig. 6 we show the boxplot of the MAEs of the predictions of PulseFormer on egoPPG-DB by split. The interquartile range across all splits is between 1.7 and 10.5 bpm.
Figure 6: Boxplot of the MAEs of the predictions of PulseFormer.
16 Cross-dataset evaluation
We evaluated PulseFormer and the two strongest baselines when training on three conventional rPPG datasets (MMPD [82], UBFC-rPPG [6], and PURE [79]) and testing on egoPPG-DB (Tab. 7), and vice versa (Tab. 8). For the rPPG datasets, we extracted the eye region using MediaPipe [47], resized to $48\times 128$, and converted to grayscale. PulseFormer consistently outperforms the baselines across all scenarios and datasets (except one case), showing strong generalization to unseen data. Please note that we can only evaluate PulseFormer w/o MITA as conventional rPPG datasets do not contain IMU data from the participants’ heads.
Train Set Model MAE MAPE
PhysFormer 20.56 27.06
MMPD FactorizePhys Not converging
PulseFormer w/o MITA 13.66 16.64
PhysFormer 18.32 23.63
UBFC-rPPG FactorizePhys 18.58 24.46
PulseFormer w/o MITA 14.83 18.57
PhysFormer 24.39 24.94
PURE FactorizePhys 13.20 15.44
PulseFormer w/o MITA 12.99 13.46
Table 7: Results (MAE) when training on conventional rPPG datasets and testing on egoPPG-DB.
Model MMPD UBFC-rPPG PURE
MAE MAPE MAE MAPE MAE MAPE
PhysFormer 11.76 14.57 16.80 16.46 23.89 37.50
FactorizePhys 12.06 15.11 14.28 14.98 26.10 40.62
PulseFormer (ours) 11.48 15.08 15.09 15.81 23.56 36.71
Table 8: Results (MAE) when training on egoPPG-DB and testing on conventional rPPG datasets.
17 HR distribution
egoPPG-DB exhibits the widest HR range (44–164 bpm, see Fig. 7) and significantly more motion (e.g., dancing, exercise bike) than other evaluated rPPG datasets, where participants typically sit calmly at a table.
Figure 7: Boxplot of HRs of egoPPG-DB and three rPPG datasets.
18 Downstream performance comparison
HR features from the other evaluated baselines perform progressively worse than those from PulseFormer when used for proficiency estimation on EgoExo4D, highlighting the importance of accurate HR estimation for downstream tasks (see Tab. 9).
Model Ego+HR Exo+HR Ego+Exo+HR
FactorizePhys 44.62 36.72 40.13
PhysFormer 44.39 36.66 43.07
PulseFormer (ours) 45.29 37.67 43.94
Table 9: Downstream performance (accuracy) on EgoExo4D using the HR predictions from the three best baseline models.
Dataset Part. Frames Hours Tasks
PURE [79] 10 110 K 1 Resting, talking, small head movements
MAHNOB-HCI [77] 27 2.6 M 12 Watching videos
MMPD [82] 33 1.2 M 11 Resting, head rotation, selfie videos
UBFC-rPPG [6] 43 150 K 1.5 Gaming on a computer
UBFC-PHYS [70] 56 2.4 M 19 Resting, Trier Social Stress Test
OBF [44] 106 3.8 M 18 Resting with varying HR levels
SCAMPS (synthetic) [54] 2800 1.7 M 16 Different facial actions
Table 10: Summary of existing datasets for rPPG.
@@ 18 Downstream performance comparison
Walking Walk back to the physical Walk back to the physical location of the start either up the stairs
location of the start or using the elevator.